# Decomposing Representation Space into Interpretable Subspaces with Unsupervised Learning

### v1

@@ Results
We apply NDM to find subspace partitions in GPT-2 Small (Radford et al., 2019), details are described in App. D.3. The results of NDM using the best hyperparameters are shown in Table 1, together with 4 baselines. As our method is unsupervised and purely computed from activations, there are no competitive baselines. Though there are a few cases where Gini coefficient is low for NDM, the overall average value shows a strong concentration of target information (0.71, significantly higher than 0.6). In App. D.5 we show ablation studies, providing experimental data on how different factors affect final results. Meanwhile, the estimated MI between subspaces decreases significantly after training (Fig. 14), matching the explanation of NDM’s training objective.
Figure 2: Preimages of the subspace activation corresponding to the same token “a” in 4 different subspaces of the residual stream $\boldsymbol{h}^{4,post}$ in GPT-2 Small. The highlighted tokens are those the activations are taken from. These activations are projected into the subspaces and the first column “sim” is the similarity between this projection in the preimage and in the query activation. The second column “pos” is the position indices of the highlighted tokens. We are showing only the most recent part of the context and as well as 5 future tokens, the actual input sequences are much longer. Subspace dimensions: $d_{2}=96,d_{6}=64,d_{7}=32,d_{8}=32$. Interpretation of encoded information: (a) current token “a”. (b) the prior token “under”. (c) current position, around 118. (d) the topic, the overall context is about licence/permission. In sum, we can see in different subspaces of the same layer, different aspects of the current context are encoded.
@@ 5.2 Qualitative Analysis of Obtained Subspaces in GPT-2
Huang et al. (2024b) show that we can interpret activations in terms of their preimage, which is the set of inputs that are mapped to the same activation by the model. Simply speaking, to interpret an activation (a vector, called query activation) at a certain place in the model (e.g., layer 3 residual stream, called activation site), we find those inputs that, when fed into the model, would give rise to roughly the same activation at that activation site. Therefore, when query activation is observed by downstream components, any input in preimage could be the current input, thus the commonality in preimage is the available information. In other words, query activation (any activation we want to interpret) “represents” the whole preimage.
@@
Instead of the whole space, we apply it to subspaces. (See discussion in App. F.1 for why it works better for subspaces). Instead of training decoders as in (Huang et al., 2024b), we collect around 1M model activations and conduct vector search among them to find preimages (see details in App. E.1). Though without a decoder the content of preimage is restricted by a fixed set of inputs (2000 sequences from MiniPile + 2000 IOI and greater-than sequences), it is easier to implement. We use cosine similarity to define the preimage, so it includes inputs whose corresponding activations are in a similar direction to the query activation. We set 0.85 as the threshold for preimages in most cases. For all qualitative results in the paper (including the APP), the subspace partition is given by the best configuration of NDM according to the test results from Sec. 5.1 (last row of Table 1).
@@ Results
In Fig. 2 we show 4 preimages corresponding to the same token at the same layer’s residual stream. We see clearly that in this contextualized representation, different aspects of the context are encoded in different subspaces, i.e., the current token, position, prior token, the topic. NDM indeed decomposes the representation space in a meaningful way. Moreover, we also find that the function/interpretation of subspaces is consistent. For example, Fig. 2(a) shows subspace 2 encodes the current token, and we find it always encodes the current token. This is also true for the other 3 subspaces shown in the other 3 subfigures (See Fig. 20 to 23 in App.). Though being consistent, we also find that the meaning of subspaces can vary, depending on factors such as whether a certain pattern occurs or a certain mechanism is triggered. We provide an extensive discussion for the qualitative results in App. E.3. In sum, qualitative results show that most of the time, the preimages are interpretable, and we can find a high-level concept to summarize the encoded information in subspaces across inputs because of good consistency. Due to space limits we only present a very limited number of examples, we strongly recommend checking more in our web application [^4] https://subspace-partition.streamlit.app.
@@ 5.3 Applicability to Larger Models
To test the applicability on larger models, we apply NDM on residual stream activations of Qwen2.5-1.5B (Qwen et al., 2025) and Gemma-2-2B (Team et al., 2024) [^5] Both without post training. As no very detailed ground-truth circuits are available for them, we design prompts that are likely to elicit separate mechanisms inside the model, and see if they are mediated by separate subspaces. If true, it is strong evidence showing our partition is meaningful.
@@
We construct inputs that elicit knowledge conflicts: “Max Planck won the Nobel Prize in Literature. A. Michael Spence won the Nobel Prize in Economics. Max Planck won the Nobel Prize in”. The model can either predict “Physics” by using its parametric knowledge, or predict “Literature” by copying the context knowledge. The first sentence associates an entity (target entity) “Max Planck” with a value (context answer) “Literature” for the field attribute, which is different from the ground truth value (parametric answer) “Physics”, followed by a correct example as part of the template, then the target entity is asked again for the same attribute. In this setting, we aim to check if there are separate subspaces functioning for either context knowledge or parametric knowledge.
We construct examples following this form with different target entities, using RAVEL dataset (Huang et al., 2024a). We do not use the dataset as the way it is designed, see App. D.6 for discussion and experimental details. To test whether a subspace plays a role in either of these two circuits quantitatively, we again use subspace patching and construct two types of counterfactual inputs (Fig. 3): 1) Param Corrupt: It aims to knock out the role of the parametric knowledge. We replace both occurrences of the target entity with a made-up entity, and other parts remain unchanged. Since the model does not have parametric knowledge for this made-up entity, the circuit for parametric knowledge is broken, while the circuit for context knowledge should not be affected much. 2) Context Corrupt: We replace only the first occurrence of the target entity with a made-up entity, other parts remain unchanged. This would presumably only break the context knowledge retrieval, as in the last sentence the target entity does not match any previous entity. We filter the data to ensure the model has parametric knowledge, resulting in 552 entities for Qwen2.5-1.5B and 719 entities for Gemma-2-2B, and construct the same number of prompts and measure the average patching effect over them. When doing patching, we patch all activations across token positions for a single subspace, since we do not know the underlying circuit, and our subspace partition is uniform across the entire sequence.
Figure 4: Subspace patching effect. Left: NDM partition on $\boldsymbol{h}^{11,mid}$ of Qwen2.5-1.5B. The model originally predicts the context answer around 60% of the time (“Clean Run”). If the entire space is patched by activations from Param Corrupt, the model always chooses the context answer (“d=1536”), and likewise for Context Corrupt. NDM finds subspaces that only mediate parametric knowledge routing (“$d_{2}=256$”, context answer is 100% under Param Corrupt while parametric answer rate remains the same as clean run under Context Corrupt) or only mediate context knowledge routing (“$d_{10}=64$”). See Fig. 18 in App. for baseline results. Right: NDM on $\boldsymbol{h}^{9,mid}$ of Gemma-2-2B. Similarly, some subspaces given by NDM have strong effect under one kind of counterfactual inputs but negligible effect under the other. See Fig. 19 in App. for baseline results.
@@ Results
We show results in Fig. 4. We can see that NDM finds some subspaces that have non-trivial patching effect under one kind of counterfactual inputs, indicating they play a role in for only one mechanism. For example, in $\boldsymbol{h}^{11,mid}$ of Qwen2.5-1.5B, subspace 2 seems to do the heavy lifting for routing parametric knowledge, while subspace 10 is crucial in routing context knowledge. Both subspaces are part of the same representation space. In contrast, baselines (Fig. 18, 19 in App.) typically show correlated patching effects under two types of counterfactual inputs. Importantly, we need to be careful when interpreting these results since the detailed circuits are unknown, see caveats in App. D.6. Moreover, we also try to qualitatively interpret some important subspaces, subspace 2 and 6 for Qwen2.5-1.5B. We found that the two subspaces indeed encode parametric knowledge while being in different stages of knowledge routing. Subspace 2 encodes local knowledge (consistently focused on the current entity/n-gram), while subspace 6 encodes knowledge propagated from elsewhere by attention layers (could be about a previous entity). See details in App. E.4.
@@ 6 Discussions
Future Directions
So far we have provided various evidence supporting our claim that we can decompose representation space into smaller and meaningful subspaces, without any supervision from human. These subspaces, being orthogonal to each other, can serve as new basic units or mediators for mechanistic interpretability. Compared to neurons, they capture the distributed nature of neural representations. Compared to sparse features, they provide more structure. Each subspace can be thought of as a group of infinitely many features sharing a certain higher-level concept, by interpreting some of them we can understand the whole group. Moreover, it enables us to understand the model by directly analyzing the model’s weight matrices. For example, we can understand an attention head by checking which subspace its query, key, and value matrices read from and which subspace it writes to (Elhage et al., 2021), and similarly for sub-vectors in MLP input and output matrices. By doing so, we build connections between subspaces across layers, resulting in subspace circuits. Single connection could be performed by multiple heads, thus such circuits reduce redundancy compared to model component circuits; single head can also perform multiple connections, thus we have a more fine-grained understanding of model components. We believe subspaces could be more “fundamental” than model components as they are designed to be as independent as possible. Moreover, subspace circuits would, like real transformers, operate in the same manner over the whole input sequence (uniform across token positions) as well as over different input sequences. Such circuits could be predecessors of source-code-like explanation of the model, since variables in RASP language (Weiss et al., 2021) can fundamentally be treated as subspaces of residual stream. To support this ambitious plan, the current version of subspace partition method might not be good enough. We believe there is plenty of room for further improvement. For example, allowing more flexible searching over dimension configurations (rather than having all dimensions as multiples of a fixed number). We also think a hierarchical structure of subspaces could also be helpful. For example, subspace A and B are very independent (e.g., current token and position), and A can be further partitioned into A1 and A2, which are less independent (e.g., the plain token and retrieved knowledge about it). More importantly, there could be potentially different ways achieving the same goal (e.g., by minimax approaches, see App. F.3). We also call for efforts from industry as this unsupervised method should benefit from more data and compute.
Limitations
We think our current approach is more like a prototype showing a promising direction rather than the optimal solution. For example, it is very likely that the model uses intermediate variables that take value from a small set (e.g., sentiment), thus encoded in tiny subspaces. But currently we start merging from subspaces of at least 32 dimension, thus only providing a coarse-grained partition. Moreover, as we interpret directions in subspaces by using cosine similarity to define preimages, we rely on the Linear Representation Hypothesis (Elhage et al., 2022; Park et al., 2023), which could be untrue in some cases (Csordás et al., 2024). Furthermore, as we mentioned in App. E.3, we do not find clear interpretations for some small subspaces. Besides reasons mentioned before, it could be because the orthogonal matrices are not trained well, or because we assume strong orthogonality between meaningful subspaces, or because some abstract variables used internally by the model are hard to interpret by only reading inputs. For example, a subspace might control which region certain heads attend to. We might need to contextualize a subspace in terms of subspaces connects to it to better explain.
@@ Appendix B Methodology Details
Activations Buffer B, $[\boldsymbol{h}_{1},\cdots,\boldsymbol{h}_{N}]$, threshold $\tau$, initial configuration $c=[d_{1},\ldots,d_{S}]$ where $\sum_{s=1}^{S}d_{s}=d$, initial orthogonal matrix $R\leftarrow I$, batch size $b$, search steps $n$, block size $m$, merging start delay $p$, merging interval $t$, maximum update steps $K$.
@@
[…]
@@
     $H_{k}$leftarrow B.next() $\triangleright$ a batch of size $m$, without being deleted from B
@@
     Sort by $\frac{{I}_{[s_{1}][s_{2}]}}{d_{s_{1}}+d_{s_{2}}}$ and select top $S/8$ non-intersecting pairs $>\tau$
@@ B.1 Total correlation Perspective
The entropy of continuous random vector can be estimated based on the distance to kth-nearest neighbor [Kozachenko and Leonenko, 1987]. Formally, given random samples $(\boldsymbol{x_{1}}\cdots\boldsymbol{x_{N}})$ of the random vector X, the entropy estimate $\hat{H}(X)$ is as follows
$\hat{H}(X)=\frac{1}{N}\sum_{n=1}^{N}\log\epsilon(n)+\text{const.}$. (6)
$\epsilon(n)$ is twice the distance from $\boldsymbol{x}_{n}$ to its kth neighbor, const. is a constant if N and k are fixed. Meanwhile, the total correlation [Watanabe, 1960] or multi-information [Studenỳ and Vejnarová, 1998] is defined as
@@ B.2 Algorithm Details
Here we supplement Sec. 3 and Algorithm 1 with some important details and explanations. The MI is measured by KSG estimator [Kraskov et al., 2004]. We use the first algorithm presented in the paper (Estimator $I^{(1)}(X,Y)$) and normalize the variance of each subspace before estimation, by dividing subspace activations by the square root of the sum of the variance of each dimension in the subspace. Note that MI is invariant under this transformation in theory but it helps reduce errors in practice [Kraskov et al., 2004]. We measure pair-wise MI for subspaces, and then divide them by the sum of dimensions of subspace pairs. We use subspace dimension as a very rough estimate of entropy, which forms the upper-bound of MI. So the intuition behind normalizing MI with dimensionality is to roughly estimate the ratio of entropy removed after the other variable is observed. We set a threshold for normalized MI, we select and sort all pairs whose normalized MI is above the threshold. We then iterate over these pairs, selecting pairs that do not intersect with selected pairs. For example, for subspace a, b, c, if (a, b) is selected, (b, c) will not be selected. We merge at most $S/8$ pairs, where $S$ is the number of subspaces. Note that merging does not change the weights of $\mathbf{R}$ (or just reordering), merging affects the training as we now search the nearest neighbor in the joint space and optimize the distance in the joint space. When we merge subspaces, the optimizer states (for the whole $\mathbf{R}$ are also re-initialized.
@@
Lastly, to clarify which hyperparameters we refer to throughout the rest of the paper. We assign them the following names: the initial dimensionality for subspaces is referred as unit size (i.e. $d_{1},\cdots d_{S}$ in Algorithm 1 which are equal at the beginning). We search the nearest neighbor among $N$ activations (including the point itself, so only $N-1$ candidates), we refer this as search number (i.e. equal to $n\times m$ in Algorithm 1). The orthogonal matrix is first trained for some steps, during this stage no merging happens, we call it merging start delay. After this stage, we measure MI and merge subspaces every $t$ step, we call $t$ merging interval. Lastly, the normalized MI threshold $\tau$ is called merging threshold.

### v2

@@ Results
We apply NDM to find subspace partitions in GPT-2 Small (Radford et al., 2019), details are described in App. D.3. The results of NDM using the best hyperparameters are shown in Table 1, together with 5 baselines, which, like NDM, are unsupervised and purely computed from activations. As we can see, NDM clearly outperform baselines. Though there are a few cases where Gini coefficient is low for NDM, the overall average value shows a strong concentration of target information (0.71, significantly higher than 0.6). Figure 22, 29 and 30 show the qualitative examples of the subspaces obtained by NDM with the highest effect in Test 1, 2, 5 respectively (See later section for how to read them). In App. D.5 we show ablation studies, providing experimental data on how different factors affect final results. Meanwhile, the estimated MI between subspaces decreases significantly after training (Fig. 15), matching the explanation of NDM’s training objective.
Figure 2: Illustration of our subspace decomposition. On the right, the interpretation of the column of token “a” is obtained by inspecting Figure 3. Likewise, each row from bottom to top corresponds to Figure 21, 23(d), 22, 24, respectively.
Figure 3: Preimages of the subspace activation corresponding to the same token “a” in 4 different subspaces of the residual stream $\boldsymbol{h}^{4,post}$ in GPT-2 Small. Tokens highlighted with red background are those where the activations are taken from. Tokens highlighted with dashed boxes reflect the encoded information. These activations are projected into the subspaces and the first column “sim” is the similarity between this projection in the preimage and in the query activation. The second column “pos” is the position indices of the highlighted tokens. We are showing only the most recent part of the context and as well as 5 future tokens, the actual input sequences are much longer. Interpretation of encoded information: (a) current token “a”. (b) the prior token “under”. (c) current position, around 118. (d) the topic, the overall context is about licence/permission. In sum, we can see in different subspaces of the same layer, different aspects of the current context are encoded.
@@ 5.2 Qualitative Analysis of Obtained Subspaces in GPT-2
(Huang et al., 2024b) show that we can interpret activations in terms of their preimage, which is the set of inputs that are mapped to the same activation by the model. Simply speaking, to interpret an activation (a vector, called query activation) at a certain place in the model (e.g., layer 3 residual stream, called activation site), we find those inputs that, when fed into the model, would give rise to roughly the same activation at that activation site. Therefore, when query activation is observed by downstream components, any input in preimage could be the current input, thus the commonality in preimage is the available information. In other words, query activation (any activation we want to interpret) “represents” the whole preimage.
@@
Instead of the whole space, we apply it to subspaces. (See discussion in App. G.1 for why it works better for subspaces). Instead of training decoders as in (Huang et al., 2024b), we collect around 1M model activations and conduct vector search among them to find preimages (see details in App. E.1). Though without a decoder the content of preimage is restricted by a fixed set of inputs (2000 sequences from MiniPile + 2000 IOI and greater-than sequences), it is easier to implement. We use cosine similarity to define the preimage, so it includes inputs whose corresponding activations are in a similar direction to the query activation. We set 0.85 as the threshold for preimages in most cases. For all qualitative results in the paper (including the web application), the subspace partition is given by the best configuration of NDM according to the test results from Sec. 5.1 (last row of Table 1).
@@ Results
Fig. 2 illustrates our findings. In Fig. 3 we show 4 preimages corresponding to the same token at the same layer’s residual stream. We see clearly that in this contextualized representation, different aspects of the context are encoded in different subspaces, i.e., the current token, position, prior token, the topic. NDM indeed decomposes the representation space in a meaningful way. Moreover, we also find that the function/interpretation of subspaces is consistent. For example, Fig. 3(a) shows subspace 2 encodes the current token, and we find it always encodes the current token. This is also true for the other 3 subspaces shown in the other 3 subfigures (See Fig. 21 to 24 in App.). Though being consistent, we also find that the meaning of subspaces can vary, depending on factors such as whether a certain pattern occurs or a certain mechanism is triggered. We provide an extensive discussion for the qualitative results in App. E.3. In sum, qualitative results show that most of the time, the preimages are interpretable, and we can find a high-level concept to summarize the encoded information in subspaces across inputs because of good consistency. Due to space limits we only present a very limited number of examples, we strongly recommend checking more in our web application [^4] https://subspace-partition.streamlit.app.
@@ 5.3 Applicability to Larger Models
To test the applicability on larger models, we apply NDM on residual stream activations of Qwen2.5-1.5B (Qwen et al., 2025) and Gemma-2-2B (Team et al., 2024) [^5] Both without post training. As no very detailed ground-truth circuits are available for them, we design prompts that are likely to elicit separate mechanisms inside the model, and see if they are mediated by separate subspaces. This would constitute strong evidence that our partition is meaningful.
@@
We construct inputs that elicit knowledge conflicts: “Max Planck won the Nobel Prize in Literature. … Max Planck won the Nobel Prize in”. The model can either predict “Physics” (parametric answer) by using its parametric knowledge, or predict “Literature” (context answer) by copying the context knowledge. We refer “Max Planck” as target entity. We construct examples following this form with different target entities, using RAVEL dataset (Huang et al., 2024a). We do not use the dataset as the way it is designed, see App. D.6 for discussion and experimental details. To test whether a subspace plays a role in either of these two circuits quantitatively, we again use subspace patching and construct two types of counterfactual inputs (Fig. 4): 1) Param Corrupt: It aims to knock out the role of the parametric knowledge. We replace both occurrences of the target entity with a made-up entity, and other parts remain unchanged. Since the model does not have parametric knowledge for this made-up entity, the circuit for parametric knowledge is broken, while the circuit for context knowledge should not be affected much. 2) Context Corrupt: We replace only the first occurrence of the target entity with a made-up entity, other parts remain unchanged. This would presumably only break the context knowledge retrieval, as in the last sentence the target entity does not match any previous entity. We filter entities to ensure the model has parametric knowledge, and construct more than 500 prompts for both models, and measure the average patching effect over them. When patching, we patch all activations across token positions for a single subspace, since we do not know the underlying circuit, and our subspace partition is uniform across the entire sequence.
Figure 5: Subspace patching effect. Left: NDM partition on $\boldsymbol{h}^{11,mid}$ of Qwen2.5-1.5B. The model originally predicts the context answer around 60% of the time (“Clean Run”). If the entire space is patched by activations from Param Corrupt, the model always chooses the context answer (“d=1536”), and likewise for Context Corrupt. NDM finds subspaces that only mediate parametric knowledge routing ($d_{2}=256$”, context answer is 100% under Param Corrupt while parametric answer rate remains the same as clean run under Context Corrupt) or only mediate context knowledge routing ($d_{10}=64$”). See Fig. 19 in App. for baseline results. Right: NDM on $\boldsymbol{h}^{9,mid}$ of Gemma-2-2B. Similarly, some subspaces given by NDM have strong effect under one kind of counterfactual inputs but negligible effect under the other. See Fig. 20 in App. for baseline results.
@@ Results
We show results in Fig. 5. We can see that NDM finds some subspaces that have non-trivial patching effect under one kind of counterfactual inputs, indicating they play a role in for only one mechanism. For example, in $\boldsymbol{h}^{11,mid}$ of Qwen2.5-1.5B, subspace 2 seems to do the heavy lifting for routing parametric knowledge, while subspace 10 is crucial in routing context knowledge. Both subspaces are part of the same representation space. In contrast, baselines (Fig. 19, 20 in App.) typically show correlated patching effects under two types of counterfactual inputs. Importantly, we need to be careful when interpreting these results since the detailed circuits are unknown, see caveats in App. D.6. Moreover, we also try to qualitatively interpret some important subspaces, subspace 2 and 6 for Qwen2.5-1.5B. We found that the two subspaces indeed encode parametric knowledge while being in different stages of knowledge routing. Subspace 2 encodes local knowledge (consistently focused on the current entity/n-gram), while subspace 6 encodes knowledge propagated from elsewhere by attention layers (could be about a previous entity). See details in App. E.4.
@@ 6 Discussions
Future Directions and Limitations
As we have shown, the resulting subspaces often align with meaningful “variables”, representing groups of infinitely many features sharing a higher-level concept. We believe they are better fundamental units to build circuits. Unlike feature circuits where connections are specific to concrete features, circuits defined over such “variables” capture relationships at a more abstract, high-level, and input-independent level, offering greater potential for algorithm-like description (see more discussion in App. G.3).
Our current approach shows a promising direction, but we expect that even stronger results could be achieved using substantial computational resources and engineering efforts from industry. With more compute, applications to larger models or learning better subspace partition on 2B models can produce more interesting subspaces (Compute usage of this paper is in App. H). More experiments can be done to explore more fine-grained partitions, instead of start merging from subspaces of at least 32 dimension. It is very likely that the model uses intermediate variables that take value from a small set (e.g., sentiment), thus encoded in tiny subspaces. Moreover, as we interpret directions in subspaces by using cosine similarity to define preimages, we rely on the Linear Representation Hypothesis (Elhage et al., 2022; Park et al., 2023), which could be untrue in some cases (Csordás et al., 2024). Furthermore, as mentioned in App. E.3, we do not find clear interpretations for some small subspaces. Besides reasons mentioned before, it could be because the orthogonal matrices require more training, or because we assume strong orthogonality between meaningful subspaces, or some abstract variables used internally by the model may be hard to interpret on the basis of inputs.
@@ Appendix B Methodology Details
Activations Buffer B, $\boldsymbol{h}_{1},\cdots,\boldsymbol{h}_{N}]$, threshold $\tau$, initial configuration $c=[d_{1},\ldots,d_{S}]$ where $\sum_{s=1}^{S}d_{s}=d$, initial orthogonal matrix $R\leftarrow I$, batch size $b$, search steps $n$, block size $m$, merging start delay $p$, merging interval $t$, maximum update steps $K$.
@@
  $H_{q}$ $\leftarrow$ B.pop($b$) $\triangleright$ a batch of activations, deleted after pop
@@
     $H_{k}$ $\leftarrow$ B.next() $\triangleright$ a batch of size $m$, without being deleted from B
@@
     Sort by $\frac{{I}_{[s_{1}][s_{2}]}}{d_{s_{1}}+d_{s_{2}}}$ and select top $S/8$ non-intersecting pairs $\tau$
@@ B.1 Total correlation Perspective
The entropy of continuous random vector can be estimated based on the distance to kth-nearest neighbor (Kozachenko and Leonenko, 1987). Formally, given random samples $(\boldsymbol{x_{1}}\cdots\boldsymbol{x_{N}})$ of the random vector X, the entropy estimate $\hat{H}(X)$ is as follows
$\hat{H}(X)=\frac{1}{N}\sum_{n=1}^{N}\log\epsilon(n)+\text{const.}$ (5)
$\epsilon(n)$ is twice the distance from $\boldsymbol{x}_{n}$ to its kth neighbor, const. is a constant if N and k are fixed. Meanwhile, the total correlation (Watanabe, 1960) or multi-information (Studenỳ and Vejnarová, 1998) is defined as
@@ B.2 Algorithm Details
Here we supplement Sec. 3 and Algorithm 1 with some important details and explanations. The MI is measured by KSG estimator (Kraskov et al., 2004). We use the first algorithm presented in the paper (Estimator $I^{(1)}(X,Y)$) and normalize the variance of each subspace before estimation, by dividing subspace activations by the square root of the sum of the variance of each dimension in the subspace. Note that MI is invariant under this transformation in theory but it helps reduce errors in practice (Kraskov et al., 2004). We measure pair-wise MI for subspaces, and then divide them by the sum of dimensions of subspace pairs. We use subspace dimension as a very rough estimate of entropy, which forms the upper-bound of MI. So the intuition behind normalizing MI with dimensionality is to roughly estimate the ratio of entropy removed after the other variable is observed. We set a threshold for normalized MI, we select and sort all pairs whose normalized MI is above the threshold. We then iterate over these pairs, selecting pairs that do not intersect with selected pairs. For example, for subspace a, b, c, if (a, b) is selected, (b, c) will not be selected. We merge at most $S/8$ pairs, where $S$ is the number of subspaces. Note that merging does not change the weights of $\mathbf{R}$ (or just reordering), merging affects the training as we now search the nearest neighbor in the joint space and optimize the distance in the joint space. When we merge subspaces, the optimizer states (for the whole $\mathbf{R}$ are also re-initialized.
@@
Lastly, to clarify which hyperparameters we refer to throughout the rest of the paper. We assign them the following names: the initial dimensionality for subspaces is referred as unit size (i.e. $d_{1},\cdots d_{S}$ in Algorithm 1 which are equal at the beginning). We search the nearest neighbor among $N$ activations (including the point itself, so only N1N-1 candidates), we refer this as search number (i.e. equal to $n\times m$ in Algorithm 1). The orthogonal matrix is first trained for some steps, during this stage no merging happens, we call it merging start delay. After this stage, we measure MI and merge subspaces every $t$ step, we call $t$ merging interval. Lastly, the normalized MI threshold $\tau$ is called merging threshold.